OpenStudy (anonymous):

lim 1-(1-2x)^(1/2)/x x->0 a) 2 b) 0 c) 1 d)The limit DNE e) None of the above

OpenStudy (anonymous):

DNE because nothing can be divided by 0

OpenStudy (anonymous):

dont delet your question , ARE YOU OK ??

OpenStudy (anonymous):

im not sure how to read the question

OpenStudy (anonymous):

is the X included in the power or what

OpenStudy (anonymous):

no its a variable

OpenStudy (anonymous):

yes but is EVERYTHING over X or just the power

OpenStudy (anonymous):

everything over x

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

as x approched 0

OpenStudy (anonymous):

Not true^^. Multiply top and bottom by 1+(1-2x)^(1/2) giving: \[\lim_{x \rightarrow 0}\frac{1-(1-2x)}{x(1+\sqrt{1-2x})}\] This reduces to: \[\lim_{x \rightarrow 0}\frac{2}{1+\sqrt{1-2x}}\] Evaluating gives you 1.

OpenStudy (anonymous):

ah ok so e then that was fast lol

OpenStudy (anonymous):

ah ok so e then that was fast lol

OpenStudy (radar):

\[1-(1-2x)^{1/2}\over x\]

OpenStudy (anonymous):

c my bad

OpenStudy (anonymous):

thnx alot

OpenStudy (radar):

did you consider the x still in the denominator???

OpenStudy (anonymous):

I SAY LIMIT IS DNE SINCE LIMIT FROM BOTH SIDES DONT AGREE

OpenStudy (anonymous):

No problem. The first post on here is utterly incorrect^^ The x is in the denominator. Calm down.

OpenStudy (anonymous):

I JUST WORKED IT OUT. LOOK AT WHAT I POSTED.

OpenStudy (anonymous):

And they do agree so stop posting all this ignorant stuff. http://www.wolframalpha.com/input/?i=limit+of+(1-sqrt(1-2x))%2Fx%2Cx%2C0

OpenStudy (anonymous):

lol i totally forgot my limits. i JUST remembered lo pitals rule

OpenStudy (anonymous):

You don't need l'hospital's rule.

OpenStudy (anonymous):

malevolence, you really hate mistakes..

OpenStudy (anonymous):

i worked it originally and got -1 but i see now my algebra was wrong

OpenStudy (anonymous):

I don't, everyone on the post is like FREAKING out over nothing. Just calm down. Jesus.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

well sorrrrry. i forgot my calc

OpenStudy (anonymous):

at least u knew it its kicking my retricelol

OpenStudy (anonymous):

haha dont worry, limits are easy once u practice them. after the AP exam, i forgot everything :D

OpenStudy (anonymous):

final is in 2 days so am doing problems but not in order

OpenStudy (anonymous):

ooooooooh. is this semester final or cumulative final over 2 semesters

OpenStudy (anonymous):

summer semester finals

OpenStudy (anonymous):

omg summer course?? daaaang gurl.

OpenStudy (anonymous):

am a guy ^_^

OpenStudy (anonymous):

o.O hehe im so sorry. daaang booy

OpenStudy (anonymous):

lol, back to work!!

OpenStudy (anonymous):

gudluk good sir.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

tyvm. dont reply to this one cause its gonn abe a chain of tyvm and ty's lol.

OpenStudy (anonymous):

and i meant YW not tyvm :)